OpenStudy (anonymous):

How do I simplify (2i^7)(3i-4)

OpenStudy (anonymous):

Use distributive property?

OpenStudy (anonymous):

Elaborate please... do the imaginary numbers affect it?

OpenStudy (anonymous):

that 3i has an exponent of one when multiply number with an exponent you don't add them unless they have like bases.., 2x3=6i^7+1 2i^7x4= 8i^7 6i^7-8i^7= now that our exponents are the same you must subtract.(above) which gives you.., -2i^7

OpenStudy (anonymous):

I appriciate your effort, but according to Wolfram Alpha that's incorrect: http://www.wolframalpha.com/input/?i=%282i^7%29%283i-4%29

OpenStudy (anonymous):

first, you have i^7, that's -i So you'll have (-2i)(3i-4) Then you'll have 6+8i (By using distributy)

OpenStudy (whpalmer4):

\[(2i^7)(3i-4) = 6i^7*i - 8i^7 = 6i^8 - 8i^7\] \[i^2=-1\]\[ i^4=-1*-1=1\] \[6i^8-8i^7=6*1*1-8*1*(-1)*i=6-8i\]

OpenStudy (anonymous):

i=i i^2=-1 i^3=-i i^4=1 i^5=i i^6=-1 i^7=-i i^8=1 And to infinity

OpenStudy (whpalmer4):

Sorry, \[6+8i\]

OpenStudy (anonymous):

oh okay i see where i went wrong thanks.., i should have solved for i after distributing..,